OpenStudy (anonymous):

What is the equation of a circle with a center at (0, 9) and a radius of 9?

OpenStudy (harsimran_hs4):

eq of a circle with center (h, k) and radius r is: (x-h)^2 + (y-k)^2 = r^2 in your ques : h = 0, k = 9 , r = 9 can you proceed now ?

OpenStudy (anonymous):

(x-0)^2 + (y-9)^2= 9^2 or 81?

OpenStudy (harsimran_hs4):

yes now you can leave it like this or open the squares

OpenStudy (anonymous):

Thank you!

OpenStudy (harsimran_hs4):

Welcome!!